OpenStudy (anonymous):

The average speed of one airplane is 200 mph faster than that of an automobile. If the automobile starts at 3 a.m. and the airplane starts from the same place at 11 a.m., they will have traveled the same number of miles by 1:00 p.m. Find the rate of the automobile.

OpenStudy (anonymous):

plane = 200 + car 2* plane = 10 * car solve the system of equations

OpenStudy (anonymous):

i got 2 and 10 from the time elapsed from when they started to 1pm

OpenStudy (shamim):

let the velocity of a car is\[v _{c}\]

OpenStudy (shamim):

so the velocity of plane is\[v _{c}+200\]

OpenStudy (shamim):

if the distance is \[x\]

OpenStudy (shamim):

anyway car travels for 10 hrs

OpenStudy (shamim):

so\[v _{c} \times 10=x\]

OpenStudy (shamim):

and plane travels for 2 hrs

OpenStudy (shamim):

so for plane\[(200+v _{c}) \times 2=x\]

OpenStudy (shamim):

and\[(200+v _{c}) \times 2=v _{c} \times 10\]

OpenStudy (shamim):

now solve it

OpenStudy (shamim):

anyway r u following me?

OpenStudy (anonymous):

what answer u got?

OpenStudy (shamim):

\[v _{c}=50 mph\]